A PERSON DRIVES 800 MILES IN A DAY AND A HALF. USE ALGEBRA TO FIND HOW LONG IT WOULD TAKE TO DRIVE 2000 MILES.
Answer by stanbon(75887)

A PERSON DRIVES 800 MILES IN A DAY AND A HALF. USE ALGEBRA TO FIND HOW LONG IT WOULD TAKE TO DRIVE 2000 MILES.
----
Use a proportion: days/distance = days/distance
------
Solve:
x/2000 = 3.5/800
---
x = 2000(3.5)/800
x = 8.75 days
=========================
Cheers,
Stan H.
==============